OpenStudy (anonymous):

2x-3y-4z=-21 -4x+2y-3z=-14 -3x-4y+2z=-10

OpenStudy (anonymous):

if you gone do it please show me how !

OpenStudy (mertsj):

\[-3x+4y+2z = -10\]\[-4x+2y-3z = -14\]\[2x-3y-4z = -21\]

OpenStudy (anonymous):

i need to find the value of x,y,z

OpenStudy (mertsj):

Multiply the first equation by 2 and add it to the third equation. That will eliminate z

OpenStudy (mertsj):

\[-6x+8y+4z = -20\] \[2x -3y-4z = -21\]

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

The result is -4x +5y = -41

OpenStudy (mertsj):

Now multiply the top equation by 3 and the second equation by 2 and add. That will eliminate z again.

OpenStudy (mertsj):

\[-9x+12y+6z = -30\]

OpenStudy (mertsj):

\[-8x+4y-6z = -28\]

OpenStudy (mertsj):

When you add those two you get: \[-17x+16y = -58\]

OpenStudy (mertsj):

So now you have two equations that contain only x and y. Do you know how to solve that system?

OpenStudy (anonymous):

yes thanks